# Accelerating Diffusion-based Super-Resolution with Dynamic Time-Spatial Sampling

### v1

@@ Abstract
Diffusion models have gained attention for their success in modeling complex distributions, achieving impressive perceptual quality in SR tasks. However, existing diffusion-based SR methods often suffer from high computational costs, requiring numerous iterative steps for training and inference. Existing acceleration techniques, such as distillation and solver optimization, are generally task-agnostic and do not fully leverage the specific characteristics of low-level tasks like super-resolution (SR). In this study, we analyze the frequency- and spatial-domain properties of diffusion-based SR methods, revealing key insights into the temporal and spatial dependencies of high-frequency signal recovery. Specifically, high-frequency details benefit from concentrated optimization during early and late diffusion iterations, while spatially textured regions demand adaptive denoising strategies. Building on these observations, we propose the Time-Spatial-aware Sampling strategy (TSS) for the acceleration of Diffusion SR without any extra training cost. TSS combines Time Dynamic Sampling (TDS), which allocates more iterations to refining textures, and Spatial Dynamic Sampling (SDS), which dynamically adjusts strategies based on image content. Extensive evaluations across multiple benchmarks demonstrate that TSS achieves state-of-the-art (SOTA) performance with significantly fewer iterations, improving MUSIQ scores by 0.2–3.0 and outperforming the current acceleration methods with only half the number of steps.
22footnotetext:    The full version is available at Arxiv.
@@ 1 Introduction
Image super-resolution (SR) [????????] aims to reconstruct high-resolution (HR) images from low-resolution (LR) inputs. Recently, diffusion models [??] have gained attention for their ability to model complex distributions, achieving notable success in SR [??????], particularly in perceptual quality [??]. Diffusion-based image super-resolution methods take two primary approaches: integrating the low-resolution image into a task-specific denoiser [??] or adapting the reverse diffusion process of pre-trained models [???]. These methods are computationally intensive, requiring 1000 steps for training and several, such as 20 (PASD [?]), 50 (SUPIR [?]), or more steps (StableSR [?]) during testing.
@@
Efforts to accelerate denoising generation focus on sampler acceleration and distillation, achieving results in 10 or fewer steps [??]. Most Diffusion SR methods [???] adopt these general strategies without considering the unique frequency characteristics of low-level vision tasks. However, in fact, recent studies like STAR [?] have highlighted the diverse recovery of diffusion-based SR across frequency domains, suggesting the potential to learn low- and high-frequency information at different training stages. Despite these insights, these works primarily focus on modifying the training process and optimization losses. Given the availability of many large-scale open source and pre-trained diffusion SR models [???], we aim to develop a tailored training-free acceleration strategy by leveraging the characteristics of Diffusion SR methods with the spatial and frequency information, seeking to enhance the performance of these existing models at minimal cost.
@@
To analyze frequency-based disparities in the denoising process, we conduct a tiny experiment, using SUPIR, one of the latest typical state-of-the-art Diffusion SR methods, on the RealPhoto60 dataset [?]. RealPhoto60 comprises 60 real-world images from common benchmarks. To explore the frequency characteristics, we applied Fourier transformation [?] to categorize spectra into low, medium, and high-frequency signals. To observe the time domain dynamics, we recorded the signal-to-noise ratio (SNR) of intermediate and final outputs over the 100-step inference. As shown in Fig. 1.a, SNR improvements were most pronounced in the later stages across all frequency bands. Notably, unlike low and medium frequency components, high-frequency components uniquely exhibited visible SNR gains in the early stages, indicating the critical role of early denoising in restoring high-frequency details.
@@
Furthermore, we analyzed spatial variations in high-frequency recovery by cropping RealPhoto60 samples into 128-sized patches and categorizing them as smooth, medium-textured, or highly textured based on variance. We recorded the SNR of high-frequency signals for each category throughout the iterations (Fig. 1.b), and analyzed the noise magnitude changes (Fig. 1.c). For an intuitive comparison, Fig. 1.d illustrates the denoising process for a typical sample with both low- and high-frequency regions, showing that smooth areas recover early, while high-frequency regions, such as fur, recover significantly in the final 0–100 steps. Spatially, regions with more high-frequency information concentrate recovery in the initial and final stages, whereas content variation influences denoising dynamics.
@@ 2.1 Diffusion Model in Super-Resolution
Due to their exceptional ability to generate high-quality images, diffusion-based super-resolution methods have garnered widespread attention. Early approaches leveraged low-resolution images as guidance by training a conditional DDPM [???], or conditionally steering a pre-trained DDPM [?] to perform super-resolution tasks. Recently, several studies have utilized pre-trained text-to-image (T2I) models, such as Stable Diffusion (SD) [??], harnessing learned priors to address super-resolution challenges and achieve high-quality image enhancements. These methods either train a ControlNet [????] or an additional encoder that encodes guidance features [??], both of which have demonstrated outstanding performance. Yet, as mentioned earlier, these methods require a large number of diffusion steps, resulting in high computational costs.
@@ 2.2 Diffusion Model Acceleration
Recent state-of-the-art super-resolution methods still require dozens or even hundreds of diffusion steps, even with acceleration techniques like DDIM [?], leading to significant time overhead. Reducing diffusion steps often degrades output quality. Various pruning [?], caching [?], and distillation [?] techniques have accelerated general T2I diffusion models while preserving generation quality. In super-resolution, ResShift [?] reduces diffusion steps by modeling a Residual Shifting Markov chain between high-resolution and low-resolution images but requires a considerable cost of 500K iterations training from scratch. Distillation approaches [??] have achieved one-step diffusion in super-resolution tasks, but their training costs remain high, and their performance is constrained by the teacher models. Therefore, given the training costs and limitations of existing methods, exploring a training-free acceleration framework for diffusion-based super-resolution is valuable and necessary.
@@ 3.1 Revisiting Diffusion in Super-Resolution
To investigate the relationship between frequency signals, spatial characteristics, and temporal steps in denoising, we conducted two demo experiments using the representative SOTA method SUPIR [?], a classic framework based on pre-trained SD [?] with fine-tuned ControlNet [?]. The experiments are conducted on the RealPhoto60 dataset containing 60 real-world images from benchmarks like RealSR [?], DRealSR [?], and web sources.
@@
We analyze the relationship between denoising stages and frequency-specific signal recovery by computing the signal-to-noise ratio (SNR) over 100-step inference using the Fourier Transform, with the final denoised result as a noise-free reference and intermediate results decoded from feature representations. As shown in Fig. 1.a, SNR increases significantly in the late stages (400–0 steps). Notably, high-frequency (HF) components, unlike low and medium frequencies, also exhibit visible SNR improvement in the early stages (1000–700 steps). To further investigate HF signals crucial for SR, Fig. 1.c presents the noise delta throughout denoising, where noise represents the residual with the final result, and delta denotes stepwise changes. Noise decreases in the early (100–700) and late (400–0) stages but rises in the middle, mirroring the SNR trend in Fig. 1.a. This suggests that, unlike low and medium frequencies, early denoising stages are also crucial for HF recovery, while intermediate stages may hinder optimization.
@@ 3.2 Time-Spatial-aware Sampling
Time Dynamics Sampling (TDS)
@@
Spatial Dynamic Sampling (SDS)
@@
Subsequently, we derive the non-uniform sampling strategy for each position based on the local smoothed variance $V_{g}$. Specifically, we assign an independent time scheduler to each position and adaptively adjust it based on local variations. The overall spatial-adaptive time scheduler $\mathbbm{S}$ is defined as the set of schedulers for each position:
$\mathbbm{S}=\{S^{\prime\prime}(a(v_{g_{i,j}}),n(v_{v_{g_{i,j}}}),T,T^{\prime})\|v_{g_{i,j}}\in V_{g}\},$ (5)
@@
$t_{spatial_{k}}=\{t_{k}\in\mathbbm{S}_{i,j}|i\in[0,H],j\in[0,W]\},$ (6)
where $\mathbbm{S}_{i,j}$ denotes the single scheduler in $\mathbbm{S}$ for position $(i,j)$. Considering computational efficiency and simplicity of implementation, we use linear functions for the projection $a()$ and $n()$, which are defined as follows:
@@ 4.1 Experiment Setup
Our approach modifies only the sampling strategy during inference, requiring no extra training. Experiments used an A800 GPU (80GB) and official codebases with pre-trained weights. PyTorch framework was used for implementation, and the detailed hyperparameters are in the full version.
@@
Performance was evaluated on both synthetic and real-world data. Synthetic data used DIV2K [?] with LR images degraded by BSRGAN [?]. Real-world evaluations used DRealSR [?], RealPhoto60 [?], and RealSR [?], along with face SR benchmarks WebPhoto-Test [?] and LFW-test [?]. RealPhoto60, WebPhoto, and LFW were tested at ×2 scale, and others at ×4.
@@
For the quantitative evaluation of super-resolution, we employed widely used perceptual quality metrics, including NIQE [?], CLIPIQA [?], MUSIQ [?], and QAlign [?], to compare the performance of the competing methods. Additionally, we provide the full-reference metrics including PSNR [?], SSIM [?], and LPIPS [?] in the full version.
@@ 4.2 Comparison with State-of-the-Art Methods
For a comprehensive comparison, we integrate the TSS framework with three recent SOTA diffusion-based super-resolution (SR) methods: StableSR [?], SUPIR [?], and PASD [?]. Their performance is evaluated against state-of-the-art (SOTA) SR approaches, including Real-ESRGAN [?], BSRGAN [?], SwinIR [?], SinSR [?], and ResShift [?], using both synthetic and real-world datasets. Real-ESRGAN, BSRGAN, and SwinIR are traditional deep learning-based SR frameworks using CNNs or transformers, while SinSR and ResShift are diffusion-accelerated SR methods leveraging distillation and fine-tuning, respectively. For fairness, the results are obtained from official codebases and pre-trained models. Quantitative evaluation and qualitative comparisons are presented in Tab. 1 and Fig. 5, respectively. As presented in Tab. 1, our SUPIRTSS outperforms the state-of-the-art acceleration stra-
@@
tegies ReShift and SinSR on most benchmarks, achieving a 1.14–5.05 improvement in NIQE [?] and a 0.09–1.52 improvement in QAlign [?] across both synthetic and real datasets. Notably, while baseline SUPIR’s 7-step QAlign results are initially inferior to ReShift and SinSR on real datasets, the incorporation of the TSS framework enables a QAlign improvement of 0.09–1.24 over SinSR and 0.12–1.58 over ReShift across all datasets. Although SinSR requires only a single iteration, its distillation strategy involves additional training costs, whereas TSS achieves comparable results without any extra training. Furthermore, compared to ReShift, which requires training and 15 inference steps, SUPIRTSS delivers superior performance across all metrics except CLIPIQA [?], achieving these results in just 7 iterations, less than half the steps. Fig. 5 provides examples of synthetic and real degradation. Taking the first example in Fig. 5 as an example, most existing methods, including CNN- and Transformer-based approaches and diffusion acceleration strategies, fail to generate the region with high-frequency textures (wrinkles, beard, and eyebrow). In contrast, SUPIRTSS produces more realistic and richer high-frequency textures, as evidenced by both qualitative visual quality and quantitative results (Fig. 5, Row. 2, Col. 4) In summary, both quantitative and qualitative analyses show that TSS significantly enhances the high-frequency restoration capabilities of recent diffusion-based SR methods, achieving state-of-the-art texture generation performance on both synthetic and real-world datasets without any additional training costs.
@@ 4.3 Ablation Study
To validate the effectiveness of Time Dynamic Sampling (TDS), we compare SUPIR with its TDS-enhanced version on the RealPhoto60 dataset. As shown in Tab. 2, TDS improves quantitative metrics, achieving a 20% increase in QAlign[?] and a 35% increase in CLIPIQA [?] at 7 steps. Fig. 6 (Col. 3 vs. 1) visually compares SUPIR results with and without TDS on real samples. TDS enhances the butterfly texture reconstruction while uniform sampling introduces artifacts like color blocking. For further analysis, Fig. 7 illustrates the relationship between high-frequency SNR and the number of denoising steps. Compared to the original SUPIR (blue), SUPIR+TDS (green) consistently achieves higher SNR at the same step count, even with fewer total steps, indicating its effectiveness in high-frequency signal restoration.
@@
Metrics Unified Later Stage Only TDS
@@
Metrics Unified Ununified
@@
Qalign $\uparrow$ 3.62 4.3305 4.2458 4.35
Table 4: Comparison of ununified functions used in Time Dynamic Sampling. The detailed function description is in the full version.
@@
Figure 8: Last 4 step in PASD’s 7-step denoising process.”+TDS” strategies fixes param A and N to their min and Max value.
Ununified Function Selection
We evaluated the impact of different non-uniform functions on sampling performance, including trigonometric, exponential, and polynomial functions. As shown in Tab. 4, all three functions outperform the baseline with a 0.6–0.75 QAlign improvement, showing that inhomogeneous sampling aligns with high-frequency signal recovery regardless of function type. For better controllability, we select the polynomial function as the final setting.
@@
To assess the necessity of Spatial Dynamic Sampling (SDS), we compare SUPIR with TDS alone and the full TSS (TDS+SDS) on RealPhoto60. As shown in Tab. 2 (Col. 2 and 3), the inclusion of SDS brings quantitative metric improvements over TDS alone. Moreover, Fig. 6 further illustrates this effect. TDS alone (Col. 3) applies a spatial-uniform scheduler, resulting in less sharp butterfly wings and an insufficiently smooth background. In contrast, SDS enables spatial-wise adaptive sampling, enhancing both texture-rich and smooth regions.

### v2

@@ Abstract
Diffusion models have gained attention for their success in modeling complex distributions, achieving impressive perceptual quality in SR tasks. However, existing diffusion-based SR methods often suffer from high computational costs, requiring numerous iterative steps for training and inference. Existing acceleration techniques, such as distillation and solver optimization, are generally task-agnostic and do not fully leverage the specific characteristics of low-level tasks like super-resolution (SR). In this study, we analyze the frequency- and spatial-domain properties of diffusion-based SR methods, revealing key insights into the temporal and spatial dependencies of high-frequency signal recovery. Specifically, high-frequency details benefit from concentrated optimization during early and late diffusion iterations, while spatially textured regions demand adaptive denoising strategies. Building on these observations, we propose the Time-Spatial-aware Sampling strategy (TSS) for the acceleration of Diffusion SR without any extra training cost. TSS combines Time Dynamic Sampling (TDS), which allocates more iterations to refining textures, and Spatial Dynamic Sampling (SDS), which dynamically adjusts strategies based on image content. Extensive evaluations across multiple benchmarks demonstrate that TSS achieves state-of-the-art (SOTA) performance with significantly fewer iterations, improving MUSIQ scores by 0.2 $\sim$ 3.0 and outperforming the current acceleration methods with only half the number of steps.
@@ 1 Introduction
Image super-resolution (SR) Wang et al. (2021b); Zhang et al. (2021); Liang et al. (2021); Qin et al. (2023); Liu et al. (2023); Zhao et al. (2023); Qin et al. (2024); Bao et al. (2025) aims to reconstruct high-resolution (HR) images from low-resolution (LR) inputs. Recently, diffusion models Ho et al. (2020); Song et al. (2020) have gained attention for their ability to model complex distributions, achieving notable success in SR Chen et al. (2024); Wang et al. (2024b); Yang et al. (2025); Yu et al. (2024); Wu et al. (2024b); Qu et al. (2025), particularly in perceptual quality Wang et al. (2023); Wu et al. (2023). Diffusion-based image super-resolution methods take two primary approaches: integrating the low-resolution image into a task-specific denoiser Saharia et al. (2022); Wang et al. (2024a) or adapting the reverse diffusion process of pre-trained models Wu et al. (2024b); Yang et al. (2025); Yu et al. (2024). These methods are computationally intensive, requiring 1000 steps for training and several, such as 20 (PASD Yang et al. (2025)), 50 (SUPIR Yu et al. (2024)), or more steps (StableSR Wang et al. (2024a)) during testing.
@@
Efforts to accelerate denoising generation focus on sampler acceleration and distillation, achieving results in 10 or fewer steps Yue et al. (2024); Wang et al. (2024b). Most Diffusion SR methods Wang et al. (2024a); Yu et al. (2024); Yang et al. (2025) adopt these general strategies without considering the unique frequency characteristics of low-level vision tasks. However, in fact, recent studies like STAR Xie et al. (2025) have highlighted the diverse recovery of diffusion-based SR across frequency domains, suggesting the potential to learn low- and high-frequency information at different training stages. Despite these insights, these works primarily focus on modifying the training process and optimization losses. Given the availability of many large-scale open source and pre-trained diffusion SR models Yang et al. (2025); Yu et al. (2024); Wang et al. (2024a), we aim to develop a tailored training-free acceleration strategy by leveraging the characteristics of Diffusion SR methods with the spatial and frequency information, seeking to enhance the performance of these existing models at minimal cost.
@@
To analyze frequency-based disparities in the denoising process, we conduct a tiny experiment, using SUPIR, one of the latest typical state-of-the-art Diffusion SR methods, on the RealPhoto60 dataset Yu et al. (2024). RealPhoto60 comprises 60 real-world images from common benchmarks. To explore the frequency characteristics, we applied Fourier transformation Cochran et al. (1967) to categorize spectra into low, medium, and high-frequency signals. To observe the time domain dynamics, we recorded the signal-to-noise ratio (SNR) of intermediate and final outputs over the 100-step inference. As shown in Fig. 1.a, SNR improvements were most pronounced in the later stages across all frequency bands. Notably, unlike low and medium frequency components, high-frequency components uniquely exhibited visible SNR gains in the early stages, indicating the critical role of early denoising in restoring high-frequency details.
@@
Furthermore, we analyzed spatial variations in high-frequency recovery by cropping RealPhoto60 samples into 128-sized patches and categorizing them as smooth, medium-textured, or highly textured based on variance. We recorded the SNR of high-frequency signals for each category throughout the iterations (Fig. 1.b), and analyzed the noise magnitude changes (Fig. 1.c). For an intuitive comparison, Fig. 1.d illustrates the denoising process for a typical sample with both low- and high-frequency regions, showing that smooth areas recover early, while high-frequency regions, such as fur, recover significantly in the final $0\sim 100$ steps. Spatially, regions with more high-frequency information concentrate recovery in the initial and final stages, whereas content variation influences denoising dynamics.
@@ 2.1 Diffusion Model in Super-Resolution
Due to their exceptional ability to generate high-quality images, diffusion-based super-resolution methods have garnered widespread attention. Early approaches leveraged low-resolution images as guidance by training a conditional DDPM Ho et al. (2020); Kawar et al. (2022); Saharia et al. (2022), or conditionally steering a pre-trained DDPM Choi et al. (2021) to perform super-resolution tasks. Recently, several studies have utilized pre-trained text-to-image (T2I) models, such as Stable Diffusion (SD) Rombach et al. (2022); Podell et al. (2023), harnessing learned priors to address super-resolution challenges and achieve high-quality image enhancements. These methods either train a ControlNet Zhang et al. (2023); Yang et al. (2025); Yu et al. (2024); Chen et al. (2024) or an additional encoder that encodes guidance features Wu et al. (2024b); Wang et al. (2024a), both of which have demonstrated outstanding performance. Yet, as mentioned earlier, these methods require a large number of diffusion steps, resulting in high computational costs.
@@ 2.2 Diffusion Model Acceleration
Recent state-of-the-art super-resolution methods still require dozens or even hundreds of diffusion steps, even with acceleration techniques like DDIM Song et al. (2020), leading to significant time overhead. Reducing diffusion steps often degrades output quality. Various pruning Zhu et al. (2024), caching Ma et al. (2024), and distillation Yin et al. (2024) techniques have accelerated general T2I diffusion models while preserving generation quality. In super-resolution, ResShift Yue et al. (2024) reduces diffusion steps by modeling a Residual Shifting Markov chain between high-resolution and low-resolution images but requires a considerable cost of 500K iterations training from scratch. Distillation approaches Wu et al. (2024a); He et al. (2024) have achieved one-step diffusion in super-resolution tasks, but their training costs remain high, and their performance is constrained by the teacher models. Therefore, given the training costs and limitations of existing methods, exploring a training-free acceleration framework for diffusion-based super-resolution is valuable and necessary.
@@ 3.1 Revisiting Diffusion in Super-Resolution
To investigate the relationship between frequency signals, spatial characteristics, and temporal steps in denoising, we conducted two demo experiments using the representative SOTA method SUPIR [2024], a classic framework based on pre-trained SD [2022] with fine-tuned ControlNet [2023]. The experiments are conducted on the RealPhoto60 dataset containing 60 real-world images from benchmarks like RealSR [2020], DRealSR [2020], and web sources.
@@
We analyze the relationship between denoising stages and frequency-specific signal recovery by computing the signal-to-noise ratio (SNR) over 100-step inference using the Fourier Transform, with the final denoised result as a noise-free reference and intermediate results decoded from feature representations. As shown in Fig. 1.a, SNR increases significantly in the late stages (400 $\sim$ 0 steps). Notably, high-frequency (HF) components, unlike low and medium frequencies, also exhibit visible SNR improvement in the early stages (1000 $\sim$ 700 steps). To further investigate HF signals crucial for SR, Fig. 1.c presents the noise delta throughout denoising, where noise represents the residual with the final result, and delta denotes stepwise changes. Noise decreases in the early (1000 $\sim$ 700) and late (400 $\sim$ 0) stages but rises in the middle, mirroring the SNR trend in Fig. 1.a. This suggests that, unlike low and medium frequencies, early denoising stages are also crucial for HF recovery, while intermediate stages may hinder optimization.
@@ 3.2 Time-Spatial-aware Sampling
3.2.1 Time Dynamics Sampling (TDS)
@@
3.2.2 Spatial Dynamic Sampling (SDS)
@@
Subsequently, we derive the non-uniform sampling strategy for each position based on the local smoothed variance $V_{g}$. Specifically, we assign an independent time scheduler to each position and adaptively adjust it based on local variations. The overall spatial-adaptive time scheduler $\mathds{S}$ is defined as the set of schedulers for each position:
$\mathds{S}=\{S^{\prime\prime}(a(v_{g_{i,j}}),n(v_{v_{g_{i,j}}}),T,T^{\prime})\|v_{g_{i,j}}\in V_{g}\},$ (5)
@@
$t_{spatial_{k}}=\{t_{k}\in\mathds{S}_{i,j}|i\in[0,H],j\in[0,W]\},$ (6)
where $\mathds{S}_{i,j}$ denotes the single scheduler in $\mathds{S}$ for position $(i,j)$. Considering computational efficiency and simplicity of implementation, we use linear functions for the projection $a()$ and $n()$, which are defined as follows:
@@ 4.1 Experiment Setup
Our approach modifies only the sampling strategy during inference, requiring no extra training. Experiments were conducted on an NVIDIA A800 GPU (80GB) using the official codebases and pre-trained weights. The PyTorch framework was used for implementation, and the detailed hyperparameters are provided in the full version.
@@
Performance was evaluated on both synthetic and real-world data. Synthetic data used DIV2K [2017] with LR images degraded by BSRGAN [2021]. Real-world evaluations used DRealSR [2020], RealPhoto60 [2024], and RealSR [2020], along with face SR benchmarks WebPhoto-Test [2021a] and LFW-test [2019]. RealPhoto60, WebPhoto, and LFW were tested at ×2 scale, and others at ×4.
@@
For the quantitative evaluation of super-resolution, we employed widely used perceptual quality metrics, including NIQE [2013], CLIPIQA [2023], MUSIQ [2021], and QAlign [2023], to compare the performance of the competing methods. Additionally, we provide the full-reference metrics including PSNR [2010], SSIM [2004], and LPIPS [2018] in the full version.
@@ 4.2 Comparison with State-of-the-Art Methods
For a comprehensive comparison, we integrate the TSS framework with three recent SOTA diffusion-based super-resolution (SR) methods: StableSR [2024a], SUPIR [2024], and PASD [2025]. Their performance is evaluated against state-of-the-art (SOTA) SR approaches, including Real-ESRGAN [2021b], BSRGAN [2021], SwinIR [2021], SinSR [2024b], and ResShift [2024], using both synthetic and real-world datasets. Real-ESRGAN, BSRGAN, and SwinIR are traditional deep learning-based SR frameworks using CNNs or transformers, while SinSR and ResShift are diffusion-accelerated SR methods leveraging distillation and fine-tuning, respectively. For fairness, the results are obtained from official codebases and pre-trained models. Quantitative evaluation and qualitative comparisons are presented in Tab. 1 and Fig. 5, respectively. As presented in Tab. 1, our SUPIRTSS outperforms the state-of-the-art acceleration stra-
@@
tegies ReShift and SinSR on most benchmarks, achieving a 1.14 - 5.05 improvement in NIQE [2013] and a 0.09 - 1.52 improvement in QAlign [2023] across both synthetic and real datasets. Notably, while baseline SUPIR’s 7-step QAlign results are initially inferior to ReShift and SinSR on real datasets, the incorporation of the TSS framework enables a QAlign improvement of 0.09 - 1.24 over SinSR and 0.12 - 1.58 over ReShift across all datasets. Although SinSR requires only a single iteration, its distillation strategy involves additional training costs, whereas TSS achieves comparable results without any extra training. Furthermore, compared to ReShift, which requires training and 15 inference steps, SUPIRTSS delivers superior performance across all metrics except CLIPIQA [2023], achieving these results in just 7 iterations, less than half the steps. Fig. 5 provides examples of synthetic and real degradation. Taking the first example in Fig. 5 as an example, most existing methods, including CNN- and Transformer-based approaches and diffusion acceleration strategies, fail to generate the region with high-frequency textures (wrinkles, beard, and eyebrow). In contrast, SUPIRTSS produces more realistic and richer high-frequency textures, as evidenced by both qualitative visual quality and quantitative results (Fig. 5, Row. 2, Col. 4) In summary, both quantitative and qualitative analyses show that TSS significantly enhances the high-frequency restoration capabilities of recent diffusion-based SR methods, achieving state-of-the-art texture generation performance on both synthetic and real-world datasets without any additional training costs.
@@ 4.3 Ablation Study
To validate the effectiveness of Time Dynamic Sampling (TDS), we compare SUPIR with its TDS-enhanced version on the RealPhoto60 dataset. As shown in Tab. 2, TDS improves quantitative metrics, achieving a 20% increase in QAlign[2023] and a 35% increase in CLIPIQA [2023] at 7 steps. Fig. 6 (Col. 3 vs. 1) visually compares SUPIR results with and without TDS on real samples. TDS enhances the butterfly texture reconstruction while uniform sampling introduces artifacts like color blocking. For further analysis, Fig. 7 illustrates the relationship between high-frequency SNR and the number of denoising steps. Compared to the original SUPIR (blue), SUPIR+TDS (green) consistently achieves higher SNR at the same step count, even with fewer total steps, indicating its effectiveness in high-frequency signal restoration.
@@
Metrics Uniform Later Stage Only TDS
@@
Metrics Uniform Non-Uniform
@@
Qalign $\uparrow$ 3.62 4.33 4.25 4.35
Table 4: Comparison of non-uniform functions used in TDS. The detailed function description is in the full version.
@@
Figure 8: Last 4 steps in PASD’s 7-step denoising process. ”+TDS” strategies fixes param $a$ and $n$ to their min and Max value.
Non-Uniform Function Selection
We evaluated the impact of different non-uniform functions on sampling performance, including trigonometric, exponential, and polynomial functions. As shown in Tab. 4, all three functions outperform the baseline with a 0.6 - 0.75 QAlign improvement, showing that non-uniform sampling aligns with high-frequency signal recovery regardless of the function type. For better controllability, we select the polynomial function as the final setting.
@@
To assess the necessity of Spatial Dynamic Sampling (SDS), we compare SUPIR with TDS alone and the full TSS (TDS+SDS) on RealPhoto60. As shown in Tab. 2 (Col. 2 and 3), the inclusion of SDS brings quantitative metric improvements over TDS alone. Moreover, Fig. 6 further illustrates this effect. TDS alone (Col. 3) applies a spatial-unified scheduler, resulting in less sharp butterfly wings and an insufficiently smooth background. In contrast, SDS enables spatial-wise adaptive sampling, enhancing both texture-rich and smooth regions.